# Cross Validation in Stochastic Analytic Continuation

### v1

@@ V.2 Heisenberg Chain with Long-Range Interactions
We now turn to a case where the exact features contained in the spectral function are unknown, the unfrustrated AFM Heisenberg chain with power-law decaying interactions:
@@
Interest in this model is rooted in its possible connection to the 2D square lattice AFM Heisenberg model, as the long-range, staggered interactions allow for symmetry breaking and true long-range order to form in this 1D quantum magnet at $T=0$, effectively increasing its dimensionality [19]. The excitation spectra of the long-range Heisenberg chain is a key ingredient in understanding the nature of its ground state, and the ability to resolve sharp features with the aforementioned constrained SAC sampling schemes may help clear up the long-standing uncertainties surrounding this model. An extensive study of this system will be published in a separate paper, where spectral functions are calculated in both the Néel ordered phase ($\alpha<\alpha_{c}=2.22(1)$), as well as in the quasi long-range ordered (QLRO) phase ($\alpha>\alpha_{c}$) [20].
@@
Figure 7: Cross validation results for the $S=1/2$ Heisenberg chain with long-range interactions ($L=256$, $\beta=256$), for $\alpha=2.0$. Panels and set up are the same as in Fig. 5.
@@
The Néel phase provides an excellent application for cross validation, because of the uncertainty surrounding the nature of the ground state excitations. In the QLRO regime, it is expected that the spectral functions contains a power-law edge, in light of the BA results discussed in Sec. V.2. A previous study on a Heisenberg chain with similar long-range interactions focused only on resolving spectral functions with sharp magnon peaks when the system is in Néel phase [19]. This study used the time-dependent Density Matrix Renormalization Group (tDMRG) method, which is known to produce artificially broadened peaks due to limits on the width of the real-time window used in the calculations. The $\delta$-function peak SAC parameterization could potentially provide a significant improvement to these results.
It is also possible that on the Néel side of the quantum phase transition (QPT) the spectral function contains a power-law edge, something that has not be yet been considered due to the difficulty of resolving this type of structure using traditional analytic continuation methods. Here, we analyze results from our auxiliary paper, Ref. [20], using cross validation to test these three SAC parameterizations for $\alpha=2.0$, in the Néel ordered regime but close to the QPT. In Ref. [20], a broader set of SAC parameterizations are tested, but here, we will only focus on the three discussed thus far.
Figure 8: Spectra corresponding to the three parameterizations in Fig. 7, unconstrained (a), single-peak with $A_{0}=0.80$ (b), and edge (c) . Panels and set up are the same as in Fig. 6
@@
Figure 7 shows the results of the cross validation procedure for this model. Here, we again use QMC data generated using the SSE method, for a system of length $L=256$ at inverse temperature $\beta=256$, and we consider the operator $O=S^{z}_{q=\pi/2}$. The number of validation data sets used in this run was $K=20$, which was chosen to limit the error level to $\sigma\sim 10^{-5}$.
The the validation $\chi^{2}$ curves in Fig. 7(b) suggest that single-peak parameterization now preforms best, followed by the edge parameterization, and then finally unconstrained sampling. This is even clearer when comparing $\chi_{\mathrm{val}}^{2}$ at fixed $\chi^{2}$, panel (c). The hierarchy of parameterizations is extremely stable over the entire annealing run, and a large gap separates the single peak parameterization from the other two. We contrast this with Fig. 5(c), where the single peak and edge $\chi_{\mathrm{val}}^{2}$ values are much closer, and only deviate as they approach their minima.
We can understand the difference in the behavior of the curves in Fig. 5(c) and Fig. 7(c) as a reflection of difference between the applicabilities of these spectral forms for each model. As discussed in Sec. V.1, the single peak parameterization preforming comparably to the edge parameterization for the plain Heisenberg chain can be explained by finite-size effects; only in the thermodynamic limit would one expect a completely smooth and monotonic spectral function. On the other hand, the long-range Heisenberg chain with $\alpha=2.0$ exhibits Néel order and will thus host coherent, spin-wave excitations, albeit with anomalous dispersion [31, 32, 20]. We then expect the spectral function to contain a dominant magnon peak, as reproduced by the single peak parameterization. In this case, the elevated $\chi_{\mathrm{val}}^{2}$ for the edge parameterization reflects that a monotonically decaying continuum does not follow this peak; the error contributed by the use of this improper parameterization is systematic. It will be very informative to preform this test for values of $\alpha>\alpha_{c}$, in the QRLO phase where a power law divergent spectral function is expected instead.
VI Discussion
@@
We have explored the use of cross validation in the SAC method, as both a confirmation of the optimal $\Theta$ criterion, as well as a tool to select the most likely spectral parameterization for a given model. In our test cases using synthetic data, we found excellent agreement between the optimal $\Theta$ criterion Eq. (11) and the location of the cross validation minimum. For the more complex of the two test cases, the double-Gaussian, the broad validation $\chi^{2}$ minimum reflected the difficulties posed by the relatively sharp features contained in this spectrum. In principle this spectrum can be resolved using SAC, but the unconstrained sampling introduces entropic distortions that produce noticeable deviations from the exact result. As a result, all sampling temperatures within the broad minimum produce acceptable representations of the exact spectrum, which we demonstrated by using a $\Theta$ value one standard deviation below the average minimum value. Interestingly, at this lower sampling temperature some features of the spectrum were actually reproduced better. Since using cross validation to fix $\Theta$ would be a very poor use of computational resources, it is very encouraging that the optimal $\Theta$ criterion agrees so well with the the results of the validation procedure.
@@
A promising, and more decisive use of cross validation, is in “model selection.” While the recent developments in the constrained SAC sampling schemes have allowed for the resolution of spectra with sharp features, the ill-posed nature of the analytic continuation problem still presents obstacles for the implementation of these new methods. When many different spectra give acceptable $\chi^{2}$ values, cross validation is perhaps the least biased way to determine which parameterizations is most likely. Our results for the $S=1/2$ AFM Heisenberg chain suggest that cross validation indeed can accomplish this task. Cross validation should be an extremely valuable tool in the study of highly correlated, many-body systems if it can be used to differentiate quantum phases of matter via their spectral features. Further tests on other systems with known spectral features will be very instructive.
@@
The test on the Heisenberg chain with long-range interactions also produced promising results, appearing to correctly identify which side of the QPT the system was on, based on the spectral features alone. While the exact features are only known with certainty in the limiting cases of $\alpha$, $\alpha\to 0$ and $\alpha\to\infty$, it is likely that these features persists throughout either side of the QPT. Further tests at multiple different values of $\alpha$ are still needed, and some will be presented in Ref. [20], but as a general diagnostic tool, cross validation shows much promise in the practice of numerical analytic continuation of QMC data.
@@
Our cross validation scheme is not only applicable to SAC, but can also be applied other numerical analytic continuation methods, such as MEM. Based on our tests, a promising application of cross validation would be in the selection of the default model used in MEM. While not exactly analogous to the SAC parameterization, there are similarities between these two user-controlled inputs, such as their effects on the shape, and thus entropy, of the output spectra. Often times, perturbation theory is used to select the MEM default model, where some parameter in the perturbative solution is optimized to select the “best” spectra [1]. Cross validation could be used to aide in this optimizations process or to compare entirely different default models, perhaps from different perturbative solutions of the model under consideration.
VII Acknowledgements

### v2

@@ VI.2 Heisenberg Chain with Long-Range Interactions
We now turn to a case where the exact features contained in the spectral function are unknown, the unfrustrated AFM Heisenberg chain with power-law decaying interactions [38, 39]:
@@
We note that the normalization of the coupling, $G$, chosen so that $\sum_{r}|J_{r}|=1$, differs slightly from that used in previous related studies. Interest in this model is rooted in its possible connection to the 2D square lattice AFM Heisenberg model, as the long-range, staggered interactions allow for symmetry breaking and true long-range order to form in this 1D quantum magnet at $T=0$, effectively increasing its dimensionality [25]. The excitation spectra of the long-range Heisenberg chain is a key ingredient in understanding the nature of its ground state, and the ability to resolve sharp features with the aforementioned constrained SAC sampling schemes may help clear up the long-standing uncertainties surrounding this model. An extensive study of this system will be published in a separate paper, where spectral functions are calculated in both the Néel ordered phase ($\alpha<\alpha_{c}=2.22(1)$), as well as in the quasi long-range ordered (QLRO) phase ($\alpha>\alpha_{c}$), on the other side of the quantum phase transition (QPT) [26].
@@
Figure 12: Cross validation results for the $S=1/2$ Heisenberg chain with long-range interactions ($L=256$, $\beta=256$), for $\alpha=1.5$ (panels (a)-(c)), 2.0 (panels (d)-(f)), and 3.0 (panels (g)-(i)). The spectral function considered here is for the operator $O=S^{z}_{q=\pi/2}$. Panels and set up are the same as in Fig. 10.
@@
The transition between the Néel and QLRO phases provides an excellent application for cross validation, because of the uncertainty surrounding the nature of the ground state excitations. Deep in the QLRO regime, it is expected that the spectral functions contains a power-law edge, in light of the BA results discussed in Sec. VI.2. While deep in the Néel phase, the spectral function should contain a dominant magnon-peak, in analogy to the 2D AFM Heisenberg model [29, 28]. A previous study on a Heisenberg chain with similar long-range interactions only resolved spectral functions with sharp magnon peaks, followed by weak features at higher energy, when the system is in Néel phase [25]. This study used the time-dependent Density Matrix Renormalization Group (tDMRG) method, which is known to produce artificially broadened peaks due to limits on the width of the real-time window used in the calculations. Due to the uncertainties introduced by this method, the nature of the higher-energy continuum was not examined in detail. The $\delta$-function peak SAC parameterization could potentially provide a significant improvement to these results, and the application of the edge parameterization could provide completely new insights into this model.
Figure 13: Comparison of the single-peak and edge parameterization spectral functions for the for the $S=1/2$ Heisenberg chain with long-range interactions, corresponding to the cross validation results shown in Fig. 10 panels (a)-(c). For the single-peak parameterization (blue) we use the leading $\delta$-function weight $A_{0}=0.85$ and $N_{c}=2,000$ continuum $\delta$-functions. For the edge parameterization (green), we use $N_{\omega}=640$ and the exponent $p=-1.00$.
Here, we use cross validation to test three SAC parameterizations (unconstrained, single-peak, and edge) in the Néel ordered regime ($\alpha=1.5$), in the Néel, but close to the QPT ($\alpha=2.0$), and on the QLRO side of the QPT ($\alpha=3.0$) for the spectral function of the operator $O=S^{z}_{q=\pi/2}$. In Ref. 26, a broader set of SAC parameterizations are tested and all momenta are examined, but here, we will only focus on this single momentum and the three parameterizations discussed thus far.
@@
Figure 12 shows the results of the cross validation procedure for this model. Here, we again used QMC data generated using the SSE method, for a system of length $L=256$ at inverse temperature $\beta=256$. The number of validation data sets used in this run was $K=20$, which was chosen to limit the error level to $\sigma=10^{-5}$. Just as before, the values of $A_{0}$ and $p$ used in these tests were identified using scans [26].
For all values of $\alpha$ considered here, the unconstrained sampling expectedly performs the worst of the three parameterizations. It is known that in both the large and small $\alpha$ limits, the spectral function contains a sharp edge feature, so it is likely that for intermediate values of $\alpha$, the spectral function also contains a low energy edge of some form. It is most instructive, however, to look at the behavior of the single-peak and edge curves as $\alpha$ is increased. For $\alpha=1.5$, the long-range Heisenberg chain exhibits Néel order and will thus host coherent, spin-wave excitations, albeit with anomalous dispersion [38, 40, 26]. We thus expect the spectral function to contain a dominant magnon peak, as was resolved using the single-peak parameterization in the 2D square lattice Heisenberg model [27]. Accordingly, the single-peak curve in Fig. 12(a) (blue) reaches the lowest validation $\chi^{2}$ minimum. The spectral functions produced using these two parameterizations for this value of $\alpha$ are shown in Fig. 13. While the two spectral edges align nearly perfectly, the distributions of the spectral weight above this lower bound differ. We attribute the observed difference in the validation $\chi^{2}$ values to this difference and conclude that the spectral function for $\alpha=1.5$ does not contain a monotonically decaying edge. We contrast that with the curves in Fig. 12(g). For $\alpha=3.0$, the ground state of the long-range Heisenberg chain is much closer to that of the standard Heisenberg chain discussed in Sec. VI.1. This is reflected in the cross validation $\chi^{2}$, where the edge parameterization instead reaches the lowest minimum value.
Finally, for $\alpha=2.0$ the validation $\chi^{2}$ does not appear to favor either of the single-peak or edge parameterizations due to the proximity to the QPT at $\alpha_{c}$. As $\alpha$ is increased from, the SAC parameterization that has the lowest validation $\chi^{2}$ switches over from the single-peak to the edge parameterization. This implies that there exists a crossover regime around the vicinity of $\alpha_{c}$ where cross validation is unable to identify a clear preference between the two spectral forms. The notion of a crossover between these two spectral forms is further supported by the smoothly increasing value of the optimal exponent $p$ as $\alpha$ is reduced. The edge parameterization spectra diverging more sharply as the systems enters deeper into the Néel ordered phase can be considered a natural transition to the expected dominant $\delta$-peak form. However, in the intermediate regime near $\alpha_{c}$, it is unlikely that the spectral function can be described purely by one of these two sharp features.
Different values of both $\alpha$ and momentum $q$ will be studied in Ref. 26, but these preliminary tests provide further support for the viability of cross validation as a tool for model selection in the analytic continuation of QMC-generated data.
VII Discussion
@@
We have explored the use of cross validation in the SAC method, as both a confirmation of the optimal $\Theta$ criterion, as well as a tool to select the most likely spectral parameterization for a given model. In our test cases using synthetic data, we found excellent agreement between the optimal $\Theta$ criterion Eq. (11) and the location of the cross validation minimum. For the more complex of the two test cases, the double Gaussian peak spectrum, the broad validation $\chi^{2}$ minimum reflected the difficulties posed by the relatively sharp features contained in this spectrum. In principle this spectrum can be resolved using SAC, but the unconstrained sampling introduces entropic distortions that produce noticeable deviations from the exact result. As a result, all sampling temperatures within the broad minimum produce acceptable representations of the exact spectrum, which we demonstrated by using a $\Theta$ value one standard deviation below the average minimum value. Interestingly, at this lower sampling temperature some features of the spectrum were actually reproduced better. Since using cross validation to fix $\Theta$ would be a very poor use of computational resources, it is very encouraging that the optimal $\Theta$ criterion agrees so well with the the results of the validation procedure.
A promising, and more decisive use of cross validation, is in model selection. The different entropy contents of the two unconstrained SAC parameterizations result in differing abilities to reproduce a given spectrum depending on its exact features. In some cases, it it advantageous to use the equal amplitudes parameterization, and in others the variable amplitudes parameterization preforms substantially better. In previous studies, this was only able to be shown retrospectively, when comparing the SAC results to an artificial spectrum used to generate synthetic data. Here, we showed that cross validation can indeed act as a tool for model selection when comparing the two different unconstrained SAC parameterizations. This will be extremely valuable in practice, as often times these two parameterizations produce considerably different results, which can be even more dramatic when other parameterizations are considered, such as the fixed grid method [20] or the orthogonal polynomial representation [41]. When there is no artificial spectrum to compare to, it is impossible to determine which of these spectra is best.
@@
Cross validation as a tool for model selection was also applied to the constrained SAC parameterizations. While the recent developments in the constrained sampling schemes have allowed for the resolution of spectra with sharp features, the ill-posed nature of the analytic continuation problem still presents obstacles for the implementation of these new methods. When many different spectra give acceptable $\chi^{2}$ values, cross validation is the least biased way to determine which parameterization is most likely, and our results for the $S=1/2$ AFM Heisenberg chain suggest that cross validation indeed can accomplish this task. Cross validation should be an extremely valuable tool in the study of highly correlated, many-body systems if it can be used to differentiate quantum phases of matter via their spectral features. Further tests on other systems with known spectral features will be very instructive.
@@
The test on the Heisenberg chain with long-range interactions also produced promising results, correctly identify which side of the QPT the system was on based on the spectral features alone. While the exact features are only known with certainty in the limiting cases of $\alpha$, $\alpha\to 0$ and $\alpha\to\infty$, it is likely that these features persists throughout either side of the QPT. Further tests at more values of $\alpha$ and $q$ are still needed, and some will be presented in Ref. 26, but as a general diagnostic tool, cross validation shows much promise in the practice of numerical analytic continuation of QMC-generated data.
@@
Our cross validation scheme is not only applicable to SAC, but can also be applied other numerical analytic continuation methods, such as MEM. Based on our tests, a promising application of cross validation would be in the selection of the default model used in MEM. While not exactly analogous to the SAC parameterization, there are similarities between these two user-controlled inputs, such as their effects on the shape of the output spectra. Often times, perturbation theory is used to select the MEM default model, where some parameter in the perturbative solution is optimized to select the “best” spectra [7]. Cross validation could be used to aide in this optimizations process or to compare entirely different default models, perhaps from different perturbative solutions of the model under consideration.
Acknowledgements.
@@
Appendix A Optimal SAC Parameters for the Heisenberg Chain
When implementing the single-peak and edge SAC parameterizations, a fixed parameter must be chosen using a simple optimization procedure. In the case of the single-peak parameterization, this parameter is the weight of the macroscopic $\delta$-peak, $A_{0}$, and in the case of the edge parameterization, this parameter is the power $p$ with which the edge asymptotically diverges with. Scans over $A_{0}$ and $p$ are preformed before cross validation to determine their optimal values.
In either case, the SAC annealing routine in run multiple times with different fixed values of $A_{0}$ or $p$, where typically the values are selected from a sensibly defined grid (not too dense or too sparse). After these scans are preformed, the $\chi^{2}$ values as a function of either $A_{0}$ or $p$ are compared at a fixed value of the sampling temperature and a minimum is located. At very low values of $\Theta$, where the sampling is dominated by noise in the QMC-generated data, the $\chi^{2}$ minimum will be very shallow, if present at all, so instead an elevated temperature is used. This temperature can be chosen based on the minimum $\chi^{2}$ reach in an unrestricted sampling SAC run on the same set of data, as detailed in Ref. 20, or simply set to a value high enough for a clear minimum to emerge (but low enough so that the $\chi^{2}$ value at the minimum is still acceptable). The exact location of the minimum depends on the sampling temperature to some extent, but in most cases this dependance is very minor. Furthermore, any value of $A_{0}$ or $p$ chosen around the $\chi^{2}$ minimum will produce nearly identical spectra.
It is crucial to have as highest possible data quality when determining the optimal values of $A_{0}$ and $p$. Therefore, we use the full set of QMC-generated data for the Heisenberg chain when preforming the parameter scans. In this case, the complete $G(\tau)$ data set has an error level $\sim~{}3~{}\times~{}10^{-6}$, nearly ten times smaller than the error level of the validation data sets.
A.1 $A_{0}$ Scan
Figure 14: Results of $A_{0}$ scan for the Heisenberg chain. Panels (a) and (b) show the $\chi^{2}$ value versus $A_{0}$ ($\Delta A_{0}=0.01$) for $q=\pi/2$ and $q=3\pi/4$, respectively.
Figure 14 shows the results of the $A_{0}$ scans used to determine the optimal weight of the macroscopic $\delta$-peak for the two momenta considered, $q=\pi/2$, panel (a), and $q=3\pi/4$, panel (b). In this case, unrestricted sampling runs (i.e. $A_{0}=0$) were first preformed to determine the minimum $\chi^{2}$ values reached at the end of each anneal. This value is then used to determine the sampling temperature for the $A_{0}>0$ runs; we use a value of $\Theta$ that is slightly elevated from that corresponding to the criterion Eq. (11) for this unrestricted sampling run [20]. At this temperature, a clear minimum is present, but the $\chi^{2}$ value is still at an acceptable value. The optimal values of $A_{0}$, as determined by the results shown in Fig. 14, are $A_{0}=0.40$ and $0.35$ for $q=\pi/2$ and $q=3\pi/4$, respectively.
Figure 15: Variable amplitudes used to produce the $\delta$-function diagram in Fig. 1(d). For an exponent $p=-0.750$, the amplitudes increase as the edge is approached. The location of the equal amplitude cross-over point, $n_{0}$, is marked.
Figure 16: Results of $p$ scan for the Heisenberg chain. Panels (a) and (c) show the $\chi^{2}$ value versus $p$ ($\Delta p=0.05$) for $q=\pi/2$ and $q=3\pi/4$, respectively and panels (b) and (d) show the corresponding values of $n_{0}$.
A.2 $p$ Scan
As discussed briefly in Sec. VI.1, and in detail in Ref. 20, the edge parameterization is able to produce a spectrum with an edge that diverges asymptotically with an arbitrary exponent $p\neq-0.5$. This is achieved by varying the amplitudes of the sampled $\delta$-functions:
$A_{i}\propto i^{c},\;\;c=2p+1$ (23)
Here, the index $i\in\{1,N_{\omega}\}$ denotes the index of the sampled $\delta$-functions (note that $p=-0.5$ corresponds to $c=0$, i.e. equal amplitudes). In both the $p=0.5$ and $p\neq-0.5$ cases, the spectrum will only exhibit the power-law divergent form very close to the edge; away from this edge, the QMC-generated data instead dictates the exact shape of the monotonically decaying tail. Therefore, it is not necessary to varying the amplitudes according to Eq. (23) for all $N_{\omega}$ of the $\delta$-functions, and the instead a cross-over to equal amplitudes is implemented at index $i=n_{0}$. The exact location of $n_{0}$ is determined by the QMC-generated data, and is thus sampled along with the other updates of the spectrum within the SAC program. Figure 15 shows an example of how the amplitudes may vary with $\omega_{i}$ for an exponent $p=-3/4$, which is what was used for the for the diagram in Fig. 1(d). The cross-over point for this example was $n_{0}/N_{\omega}=0.50$, meaning that half of the amplitudes vary according to Eq. (23).
Because of this, one has to be slightly more careful when analyzing the results of a scan over $p$. If the $p$ value that minimizes the $\chi^{2}$ value is accompanied by a very small value of $n_{0}$, this implies that this optimal exponent is “fictitious.” In this case, the $\delta$-functions transition very quickly to the equal amplitude form and the asymptotic divergence $(\omega-\omega_{0})^{p}$ is only realized is a very small region of the spectrum. In contrast, if the optimal exponent has a $n_{0}$ value that is close to one, the amplitude form is applied to the majority of the $\delta$-functions and is thus realized for a large portion of the spectrum. This suggests that the accompanying low $\chi^{2}$ value indeed indicates a good statistical fit for a spectrum with this specific divergent edge form.
There are no definite rules for determining the optimal exponent under these considerations, but Fig. 16 provides two characteristic examples of how one may use the $\chi^{2}$ and $n_{0}$ values from a scan over $p$ to determine the optimal exponent. Panels (a) and (b) show the $\chi^{2}$ and $n_{0}$ values versus $-p$ for $S(q=\pi/2,\omega)$. In this case, the $\chi^{2}$ value is nearly constant for $-p<0.50$ and rapidly increases when $-p$ exceeds 0.50. The value of $n_{0}$ is small (less than $\sim 0.30$ and as small as 0.02 for $-p=0.40$) during this constant $\chi^{2}$ platform, which explains this behavior. The low values of $n_{0}$ imply that the edge is, for the most part, of the form $(\omega-\omega_{0})^{-1/2}$, which is corroborated by the fact that the $\chi^{2}$ value when $p$ is exactly $-0.50$ is roughly equal to the $\chi^{2}$ values for $-p<0.50$.
When $-p$ exceeds 0.50, the value of $n_{0}$ jumps close to one. Since the $\chi^{2}$ value also begins to increase in this region, $-p>0.50$ can be ruled out as optimal exponent values. The combination of these two trends strongly imply that $p=-0.50$ is the true exponent, which is to be fully expected for dynamic spin structure factor of the Heisenberg chain, calculated using BA solution [30, 31, 32].
Panels (c) and (d) show the results of the $p$ scan for $S(q=3\pi/4,\omega)$. Just as for $q=\pi/2$, $n_{0}$ is small for $-p<0.50$ (less than $\sim 0.20$ and as small as 0.02 for $-p=0.30$). However, in contrast to $q=\pi/2$, this range of $p$ values does not coincide with the $\chi^{2}$ minimum. Therefore, we can say with certainty that the optimal exponent does not lie within this range. In this case, there is a clear $\chi^{2}$ minimum at $-p=0.70$, but the accompanying value of $n_{0}$ is small, $\sim 0.20$. The $n_{0}$ value for $-p=0.60$ is much larger, around $0.60$, and the $\chi^{2}$ value is within 4 % of the overall minimum value. This difference in the $\chi^{2}$ is likely too small to say with certainty that $p=-0.70$ is really preferable over $p=-0.60$, so considering the much larger $n_{0}$ value, we conclude that the true optimal exponent for $q=3\pi/4$ is $p=-0.60$. This is further supported by the very good agreement between the SAC spectra with $p=-0.60$ and the BA solution, shown in Fig. 11(f).